# Attention layers provably solve single-location regression

### v1

@@ 4. Risk of the oracle and of the linear predictors
This assumption is made everywhere in the paper, even though it is not reminded explicitly at each result. This assumption is relatively mild in a high-dimensional setting where any two independent vectors uniformly distributed on the sphere are close to being orthogonal.
Oracle predictor.  Our first result characterizes the risk of the proposed transformer model (3) with oracle parameters $(k^{\star},v^{\star})$. All the proofs of the paper are deferred to the Appendix.
@@
where $\kappa:=k^{\top}k^{\star}$, $\nu:=v^{\top}v^{\star}$, $\theta:=v^{\top}k^{\star}$, $\eta:=k^{\top}v^{\star}$, and $\rho:=k^{\top}v$. A closed-form expression of $\mathcal{R}_{\lambda}^{<}$ is given in Appendix B. In particular,
@@
Asymptotic Bayes optimality.  Let us start by observing that the Bayes risk associated with problem ($P_{\mathrm{learn}}$) is larger than $\varepsilon^{2}$, which follows from elementary properties of the conditional expectation (Le Gall, 2022, Chapter 11). Indeed, using the Pythagorean theorem, one easily shows that
@@
With the scaling $\lambda\sqrt{d}\to\infty$, the argument of the first $\mathrm{erf}$ nonlinearity diverges to infinity with $d$. Thus it reaches the saturating part of $\mathrm{erf}$, so the first term in (9) converges to $\mathbb{E}[Y|\mathbb{X},J_{0}]$. On the other hand, the argument of the $\mathrm{erf}$ nonlinearities inside the sum are of order $\lambda=o(1)$. Thus they are in the linear part of $\mathrm{erf}$. Therefore, the sum consists of $L-1$ independent terms, each of magnitude $\lambda$. As a consequence, the whole sum is of order $\Theta(\lambda\sqrt{L})$, and we get
@@
Due to the scaling $\lambda\sqrt{L}\to 0$, we obtain that the second term decays to zero, and the oracle predictor therefore implements the conditional expectation of $Y$ given $\mathbb{X}$ and $J_{0}$. This is the best that we can hope for: the predictor succeeds in inferring the latent variable $J_{0}$, then gives the best possible prediction of $Y$ given $\mathbb{X}$ and $J_{0}$. We also see here the crucial role played by the nonlinearity $\mathrm{erf}$, whose linear part acts for $j\neq J_{0}$ and saturating part for $j=J_{0}$. In particular, such a reasoning would not hold if we had simply taken a linear function instead of $\mathrm{erf}$.
Linear model.  The asymptotic optimality of our oracle predictor is particularly striking in comparison to the risk of the optimal linear predictor. More precisely, let
@@
be the optimal linear predictor for the regression task ($P_{\mathrm{learn}}$). Its associated risk is $\mathcal{R}(\beta^{\star})=\mathbb{E}\big{[}(Y-(X_{1}^{\top},\ldots,X_{L}^{\top})\beta^{\star})^{2}\big{]}$. Both the optimal predictor and its risk can be explicitly characterized as follows.
@@
Let $p_{j}=\mathbb{P}(J_{0}=j)$ for $j\in\{1,\dots,L\}$. Then the optimal linear predictor is parameterized by
$\beta^{\star}=\begin{pmatrix}b_{1}v^{\star}\\ \vdots\\ b_{L}v^{\star}\end{pmatrix}\,,\quad b_{j}=\frac{\gamma^{2}p_{j}}{1+p_{j}(\gamma^{2}-1)}\,,$
and its risk is
@@
This result calls for a few comments. If the number of tokens is $L=1$ or if $J_{0}$ is a constant location (meaning that one $p_{j}$ is equal to $1$ while the others are equal to 00), then the learning problem ($P_{\mathrm{learn}}$) corresponds to a standard linear regression. In this case, $\mathcal{R}(\beta^{\star})=\varepsilon^{2}$, and the linear predictor $(X_{1},\ldots,X_{L})\mapsto(X_{1}^{\top},\ldots,X_{L}^{\top})\beta^{\star}$ achieves the Bayes risk.
Figure 3. Risk of the oracle predictor (Theorem 1, solid lines) and of the best linear predictor (Proposition 3, dashed lines), depending on the dimensions $d$ and $L$. The oracle predictor outperforms the linear predictor when scaling $d$. We take $\varepsilon^{2}=0$, $\gamma=1/\sqrt{2}$, $\lambda=1/d^{0.4}$, and all $p_{j}$ equal to $1/L$.
At the other end of the spectrum, in the case where $J_{0}$ is uniform over $\{1,\dots,L\}$, the formula for the risk of the linear predictor simplifies to $\mathcal{R}(\beta^{\star})=\varepsilon^{2}+\gamma^{2}-\frac{\gamma^{4}}{\gamma^{2}+L-1}$. When $L\to\infty$, this risk tends to $\varepsilon^{2}+\gamma^{2}$, that is, the performance of the null predictor. In other words, the optimal linear predictor performs no better than always predicting zero. More generally, this conclusion is true in any limit where $L\to\infty$ and $\max p_{j}\to 0$. This can be explained by the fact that the location of the relevant token for prediction is random, varying from sentence to sentence. Unable to leverage this latent information, the linear regressor balances all its coefficients, resulting in uniformly poor prediction performance. This stands in sharp contrast to Corollary 2, which shows that the oracle predictor $T_{\lambda}^{(k^{\star},v^{\star})}$ is able to account for the complexity of the task, at least asymptotically. This is also illustrated by Figure 3, which compares the value of the risks given by Theorem 1 and Proposition 3. However, implementing the oracle predictor requires knowledge of the parameters $k^{\star}$ and $v^{\star}$. Our goal in the next section is therefore to show that gradient descent is able to recover these parameters.
5. Gradient descent provably recovers the oracle predictor
@@
We emphasize that ${\mathcal{R}}_{\lambda}(k,v)$ is a theoretical risk, which depends on the distribution of the pair $(\mathbb{X},Y)$ (defined in Section 2). In practice, an empirical version of this risk is minimized. As we show experimentally (see Figure 5), the stochastic dynamics induced by the empirical version of the risk are qualitatively similar to the deterministic dynamics of the theoretical risk. Therefore, in the remainder of the article, we focus on the theoretical risk to avoid overcomplicating the analysis. The transition to empirical risk is left for future research.
@@
$\begin{split}k_{t+1}&=\mathrm{Proj}_{\mathbb{S}^{d-1}}(k_{t}-\alpha(I_{d}-k_{t}k_{t}^{\top})\nabla_{k}\mathcal{R}_{\lambda_{t}}(k_{t},v_{t}))=\frac{k_{t}-\alpha(I_{d}-k_{t}k_{t}^{\top})\nabla_{k}\mathcal{R}_{\lambda_{t}}(k_{t},v_{t})}{\left\|k_{t}-\alpha(I_{d}-k_{t}k_{t}^{\top})\nabla_{k}\mathcal{R}_{\lambda_{t}}(k_{t},v_{t})\right\|_{2}}\,,\\ v_{t+1}&=\mathrm{Proj}_{\mathbb{S}^{d-1}}(v_{t}-\alpha(I_{d}-v_{t}v_{t}^{\top})\nabla_{v}\mathcal{R}_{\lambda_{t}}(k_{t},v_{t}))=\frac{v_{t}-\alpha(I_{d}-v_{t}v_{t}^{\top})\nabla_{v}\mathcal{R}_{\lambda_{t}}(k_{t},v_{t})}{\left\|v_{t}-\alpha(I_{d}-v_{t}v_{t}^{\top})\nabla_{v}\mathcal{R}_{\lambda_{t}}(k_{t},v_{t})\right\|_{2}}\,,\end{split}$ (10)
where $\mathrm{Proj}_{\mathbb{S}^{d-1}}$ denotes the Euclidean projection on the unit sphere of ${\mathbb{R}}^{d}$.
@@
Experimentally, we observe in Figure 4(a) that PGD is able to recover the oracle parameters $(k^{\star},v^{\star})$. Note that running the PGD iterates (10) involves computing the gradients $\nabla_{k}\mathcal{R}_{\lambda_{t}}(k_{t},v_{t})$ and $\nabla_{v}\mathcal{R}_{\lambda_{t}}(k_{t},v_{t})$, which is non-trivial a priori. A direct approach using Monte Carlo simulations would require a large number of sample points to reduce variance, which is computationally intractable in particular in high-dimension, and in any case gives an approximate result. Instead, we leverage our closed form formula for $\mathcal{R}_{\lambda}^{<}$ from Theorem 1 to get (nearly) exact values for the gradients. Interestingly, we also observe in Figure 4(a) that $v$ aligns with $v^{\star}$ much faster than $k$ aligns with $k^{\star}$. This is typical of two-timescale dynamics, which is a common framework in analysis of non-convex learning dynamics (Heusel et al., 2017; Dagréou et al., 2022; Hong et al., 2023; Marion and Berthier, 2023; Berthier et al., 2024; Marion et al., 2024).
@@
Figure 4. Convergence of PGD to the oracle parameters. Left: Excess risk as a function of the number of steps. Middle left: Alignment $|\kappa|=|k^{\top}k^{\star}|$ and $|\nu|=|v^{\top}v^{\star}|$ with the oracle parameters. Middle right: Trajectories of $\kappa$ and $\nu$ in two repetitions of the experiments. Each repetition corresponds to a color, the trajectory starts in the middle and ends at a corner of the plot. Right: Distance to the invariant manifold ${\mathcal{M}}$. In all plots except the middle right ones, the experiment is repeated $30$ times with independent random initializations, and $95\%$ percentile intervals are plotted (but are not visible when the variance is too small). Parameters are $d=400$, $L=10$, $\gamma=\sqrt{1/2}$, and (a) $\lambda_{t}=1/(1+10^{-4}t)$, (b) $\lambda_{t}=0.1$. More details are given in Appendix D.
@@
We introduce this manifold on the one hand owing to the observation in Figure 4(a)(right) that the dynamics converge to this manifold even when initialized on the sphere, and on the other hand because this allows to reduce the problem to a lower-dimensional subspace and to simplify the expression of the risk. Cleary, due to Assumption 1, the oracle parameters $(k^{\star},v^{\star})$ belong to ${\mathcal{M}}$. A first key property of this manifold is invariance under the PGD dynamics.
@@
Take a constant inverse temperature $\lambda_{t}\equiv\lambda>0$. Then there exists $\overline{\alpha}>0$ such that, for any step size $\alpha\leqslant\overline{\alpha}$, and for a generic initialization $(k_{0},v_{0})\in\mathcal{M}$,
$(k_{t},v_{t})\xrightarrow[t\to\infty]{}\pm(k^{\star},v^{\star}).$
@@
We emphasize that the manifold ${\mathcal{M}}$ depends on the unknown parameters $k^{\star}$ and $v^{\star}$, making it impractical to initialize directly on the manifold. If the initialization is not on ${\mathcal{M}}$, more diverse phenomena are possible. As already pointed out in Figure 4(a), it is possible to obtain recovery of $(k^{\star},v^{\star})$ and convergence to the manifold ${\mathcal{M}}$ from a general initialization on the sphere. This suggests that our analysis on the manifold is relevant, and completing the analysis for a general initialization is left for future work. However, we note that using a decreasing inverse temperature sequence $\lambda_{t}$ is crucial for the recovery of $(k^{\star},v^{\star})$ when initialized out of ${\mathcal{M}}$. Indeed, to the best of our experiments, an iteration-independent choice of $\lambda$ does not consistently lead to the recovery of $k^{\star}$ and $v^{\star}$ in this case (see Appendix D). This contrasts with the behavior on the manifold proven in Theorem 5.
@@
The proof of the theorem relies on a detailed analysis of the dynamics of the PGD algorithm on the invariant manifold ${\mathcal{M}}$, in particular the properties of its stationary points. These arguments, which lie at the intersection of dynamical systems and topology, are of independent interest. A key idea is to reduce the problem to a two-dimensional system depending only on $|\kappa|=|k^{\top}k^{\star}|$ and $|\nu|=|v^{\top}v^{\star}|$. The main steps of the proof and intermediate results are given in the next section, while the core of the proofs is in Appendix A.
@@
Figure 5. Convergence of stochastic PGD to the oracle parameters from a random initialization on $(\mathbb{S}^{d-1})^{2}$. Left: Excess risk as a function of the number of steps. Middle left: Alignment $|\kappa|=|k^{\top}k^{\star}|$ and $|\nu|=|v^{\top}v^{\star}|$ with the oracle parameters. Middle right: Trajectories of $\kappa$ and $\nu$ in two repetitions of the experiment. Each repetition corresponds to a color, the trajectory starts in the middle and ends at a corner of the plot. Right: Distance to the invariant manifold ${\mathcal{M}}$. In all plots except the middle right one, the experiment is repeated $30$ times with independent random initializations, and $95\%$ percentile intervals are plotted. Parameters are $d=80$, $L=10$, $\gamma=\sqrt{1/2}$, $\lambda_{t}=2/(1+10^{-4}t)$, and a batch size of $5$. More details are given in Appendix D.
6. Outline of the proof of Theorem 5
This section outlines the essential steps for the proof of Theorem 5. For clarity, the proofs are to be found in Appendix A, except the proof of Proposition 10.

### v2

@@ 4 Risk of the oracle and of the linear predictors
This assumption is made everywhere in the remainder of the paper, even though it is not reminded explicitly at each result. It is a relatively mild assumption in a high-dimensional setting where any two independent vectors uniformly distributed on the sphere are close to being orthogonal.
Oracle predictor.
Our first result characterizes the risk of the proposed transformer model (3) with oracle parameters $(k^{\star},v^{\star})$. All the proofs of the paper are deferred to the Appendix.
@@
where $\kappa:=k^{\top}k^{\star}$, $\nu:=v^{\top}v^{\star}$, $\theta:=v^{\top}k^{\star}$, $\eta:=k^{\top}v^{\star}$, and $\rho:=k^{\top}v$. A closed-form expression of $\mathcal{R}_{\lambda}^{<}$ is given in Appendix C. In particular,
@@
Asymptotic Bayes optimality.
Let us start by observing that the Bayes risk associated with problem ($P_{\mathrm{learn}}$) is larger than $\varepsilon^{2}$, which follows from elementary properties of the conditional expectation (Le Gall, 2022, Chapter 11). Indeed, using the Pythagorean theorem, one easily shows that
@@
With the scaling $\lambda\sqrt{d}\to\infty$, the argument of the first $\mathrm{erf}$ nonlinearity diverges to infinity with $d$. Thus it reaches the saturating part of $\mathrm{erf}$, so the first term in (9) converges to $\mathbb{E}[Y|\mathbb{X},J_{0}]$. On the other hand, the argument of the $\mathrm{erf}$ nonlinearities inside the sum are of order $\lambda=o(1)$. Thus they are in the linear part of $\mathrm{erf}$. Therefore, the sum consists of $L-1$ independent terms, each of magnitude $\lambda$. As a consequence, by the central limit theorem, the whole sum is of order $\Theta(\lambda\sqrt{L})$, and we get
@@
Due to the scaling $\lambda\sqrt{L}\to 0$, the second term decays to zero, and the oracle predictor implements the conditional expectation of $Y$ given $\mathbb{X}$ and $J_{0}$. This is the best that we can hope for: the predictor succeeds in inferring the latent variable $J_{0}$, then gives the best possible prediction of $Y$ given $\mathbb{X}$ and $J_{0}$. We also see the crucial role played by the nonlinearity of $\mathrm{erf}$, whose linear part acts for $j\neq J_{0}$ and saturating part for $j=J_{0}$. In particular, the reasoning would not hold for a linear activation.
Linear model.
The asymptotic optimality of our oracle predictor is particularly striking in comparison to the risk of the optimal linear predictor. More precisely, let
@@
be the optimal linear predictor for the regression task ($P_{\mathrm{learn}}$). Its associated risk is $\mathcal{R}(\beta^{\star})=\mathbb{E}\big{[}(Y-(X_{1}^{\top},\ldots,X_{L}^{\top})\beta^{\star})^{2}\big{]}$. Both the optimal predictor and its risk can be made explicit as follows.
@@
Let $p_{j}=\mathbb{P}(J_{0}=j)$ for $j\in\{1,\dots,L\}$. Then the optimal linear predictor is parameterized by $\beta^{\star}=(b_{1}v^{\star},\dots,b_{L}v^{\star})$, with $b_{j}=\frac{\gamma^{2}p_{j}}{1+p_{j}(\gamma^{2}-1)}$, and its risk is
@@
[…]
Figure 3: Risk of the oracle predictor (Theorem 1, solid lines) and of the best linear predictor (Proposition 3, dashed lines), depending on the dimensions $d$ and $L$. The oracle predictor outperforms the linear predictor when scaling $d$. We take $\varepsilon^{2}=0$, $\gamma=1/\sqrt{2}$, $\lambda=1/d^{0.4}$, and all $p_{j}$ equal to $1/L$.
At the other end of the spectrum, in the case where $J_{0}$ is uniform over $\{1,\dots,L\}$, the formula for the risk of the linear predictor simplifies to $\mathcal{R}(\beta^{\star})=\varepsilon^{2}+\gamma^{2}-\frac{\gamma^{4}}{\gamma^{2}+L-1}$. When $L\to\infty$, this risk tends to $\varepsilon^{2}+\gamma^{2}$, that is, the performance of the null predictor. In other words, the optimal linear predictor performs no better than always predicting zero. More generally, this conclusion is true in any limit where $L\to\infty$ and $\max p_{j}\to 0$. This can be explained by the fact that the location of the relevant token for prediction is random, varying from sentence to sentence. Unable to leverage this latent information, the linear regressor balances all its coefficients, resulting in poor prediction performance. This stands in sharp contrast to Corollary 2, which shows that the oracle predictor $T_{\lambda}^{(k^{\star},v^{\star})}$ is able to account for the complexity of the task, at least asymptotically. This is also illustrated by Figure 3, which compares the value of the risks given by Theorem 1 and Proposition 3.
Naturally, implementing the attention-based oracle predictor $T_{\lambda}^{(k^{\star},v^{\star})}$ requires knowledge of the parameters $k^{\star}$ and $v^{\star}$. Our goal in the next section is therefore to show that gradient descent is able to recover these parameters.
5 Gradient descent provably recovers the oracle predictor
@@
We emphasize that ${\mathcal{R}}_{\lambda}(k,v)$ is a theoretical risk, which depends on the distribution of the pair $(\mathbb{X},Y)$ (defined in Section 2). In practice, an empirical version of this risk is minimized. As we show experimentally (see Figure 5), the stochastic dynamics induced by the empirical version of the risk are qualitatively similar to the deterministic dynamics of the theoretical risk. In the remainder of the article, we focus on the theoretical risk for simplicity, and leave the empirical risk for future research.
@@
$\begin{split}k_{t+1}&=\mathrm{Proj}_{\mathbb{S}^{d-1}}(k_{t}-\alpha(I_{d}-k_{t}k_{t}^{\top})\nabla_{k}\mathcal{R}_{\lambda_{t}}(k_{t},v_{t}))\,,\\ v_{t+1}&=\mathrm{Proj}_{\mathbb{S}^{d-1}}(v_{t}-\alpha(I_{d}-v_{t}v_{t}^{\top})\nabla_{v}\mathcal{R}_{\lambda_{t}}(k_{t},v_{t}))\,,\end{split}$ (10)
where $\mathrm{Proj}_{\mathbb{S}^{d-1}}:x\mapsto x/\|x\|_{2}$ denotes the Euclidean projection on the unit sphere of ${\mathbb{R}}^{d}$.
@@
Experimentally, we observe in Figure 4(a) that PGD is able to recover the oracle parameters $(k^{\star},v^{\star})$. Note that running the PGD iterates (10) involves computing the gradients $\nabla_{k}\mathcal{R}_{\lambda_{t}}(k_{t},v_{t})$ and $\nabla_{v}\mathcal{R}_{\lambda_{t}}(k_{t},v_{t})$, which is non-trivial a priori. A direct approach using Monte Carlo simulations would require a large number of sample points to reduce variance, which is computationally intractable in particular in high-dimension, and in any case gives an approximate result. Instead, we leverage our closed form formula for $\mathcal{R}_{\lambda}^{<}$ from Theorem 1 to get exact values for the gradients (up to numerical errors). Interestingly, we also observe in Figure 4(a) that $v$ aligns with $v^{\star}$ much faster than $k$ aligns with $k^{\star}$. This is typical of two-timescale dynamics, which is a common framework in analysis of non-convex learning dynamics (Heusel et al., 2017; Dagréou et al., 2022; Hong et al., 2023; Marion & Berthier, 2023; Berthier et al., 2024; Marion et al., 2024).
@@
Figure 4: Convergence of PGD to the oracle parameters. Left: Excess risk as a function of the number of steps. Middle left: Alignment $|\kappa|=|k^{\top}k^{\star}|$ and $|\nu|=|v^{\top}v^{\star}|$ with the oracle parameters. Middle right: Trajectories of $\kappa$ and $\nu$ in two repetitions of the experiments. Each repetition corresponds to a color, the trajectory starts in the middle and ends at a corner of the plot. Right: Distance to the invariant manifold ${\mathcal{M}}$. In all plots except the middle right ones, the experiment is repeated $30$ times with independent random initializations, and $95\%$ percentile intervals are plotted (but are not visible when the variance is too small). Parameters are $d=400$, $L=10$, $\gamma=\sqrt{1/2}$, and (a) $\lambda_{t}=1/(1+10^{-4}t)$, (b) $\lambda_{t}=0.1$. More details are given in Appendix E.
@@
We introduce this manifold on the one hand owing to the observation in Figure 4(a) (right) that the dynamics converge to this manifold even when initialized on the sphere, and on the other hand because this allows to reduce the problem to a lower-dimensional subspace and to simplify the expression of the risk. Clearly, due to Assumption 1, the oracle parameters $(k^{\star},v^{\star})$ belong to ${\mathcal{M}}$. A first key property of this manifold is invariance under the PGD dynamics.
@@
Take a constant inverse temperature $\lambda_{t}\equiv\lambda>0$. Then there exists $\overline{\alpha}>0$ such that, for any step size $\alpha\leqslant\overline{\alpha}$, and for a generic initialization $(k_{0},v_{0})\in\mathcal{M}$, $(k_{t},v_{t})\xrightarrow[t\to\infty]{}\pm(k^{\star},v^{\star}).$
@@
We emphasize that the manifold ${\mathcal{M}}$ depends on the unknown parameters $k^{\star}$ and $v^{\star}$, making it impractical to initialize directly on the manifold. If the initialization is not on ${\mathcal{M}}$, more diverse phenomena are possible. As already pointed out in Figure 4(a), it is possible to obtain recovery of $(k^{\star},v^{\star})$ and convergence to the manifold ${\mathcal{M}}$ from a general initialization on the sphere. This suggests that our analysis on the manifold is relevant, and completing the analysis for a general initialization is left for future work. However, we note that using a decreasing inverse temperature sequence $\lambda_{t}$ is crucial for the recovery of $(k^{\star},v^{\star})$ when initialized out of ${\mathcal{M}}$. Indeed, to the best of our experiments, an iteration-independent choice of $\lambda$ does not consistently lead to the recovery of $k^{\star}$ and $v^{\star}$ in this case (see Appendix E). This contrasts with the dynamics on the manifold proven in Theorem 5.
@@
The proof of the theorem relies on a detailed analysis of the dynamics of the PGD algorithm on the invariant manifold ${\mathcal{M}}$, in particular the properties of its stationary points. These arguments, which lie at the intersection of dynamical systems and topology, are of independent interest. A key idea is to reduce the problem to a two-dimensional system depending only on $\kappa=k^{\top}k^{\star}$ and $\nu=v^{\top}v^{\star}$.
Finally, numerical experiments show that a full Transformer layer is able to solve the single-location regression task. Similarly to our simplified predictor, the weights align with the oracle parameters $k^{\star}$ and $v^{\star}$. This supports the connection drawn in Section 3 between our predictor and attention layers. We refer to Appendix E for details and plots, as well as experiments on multiple-location regression, a variant of single-location regression where the output depends on several tokens.
@@
Figure 5: Convergence of online stochastic PGD to the oracle parameters from a random initialization on $(\mathbb{S}^{d-1})^{2}$. Left: Excess risk as a function of the number of steps. Middle left: Alignment $|\kappa|=|k^{\top}k^{\star}|$ and $|\nu|=|v^{\top}v^{\star}|$ with the oracle parameters. Middle right: Trajectories of $\kappa$ and $\nu$ in two repetitions of the experiment. Each repetition corresponds to a color, the trajectory starts in the middle and ends at a corner of the plot. Right: Distance to the invariant manifold ${\mathcal{M}}$. In all plots except the middle right one, the experiment is repeated $30$ times with independent random initializations, and $95\%$ percentile intervals are plotted. Parameters are $d=80$, $L=10$, $\gamma=\sqrt{1/2}$, $\lambda_{t}=2/(1+10^{-4}t)$, and a batch size of $5$. More details are given in Appendix E.
6 Conclusion
This paper introduced single-location regression, a novel statistical task where the relevant information in the input sequence is supported by a single token. We analyzed the statistical properties and optimization dynamics of a natural estimator for this task, which can be seen as a basic attention layer. We hope this work encourages further research into how Transformer architectures address sparsity and long-range dependencies, while simultaneously constructing internal linear representations of their input—–an aspect with significant implications for interpretability. Beyond NLP, potential applications include problems connected to sparse sequential modeling such as anomaly detection in time series. A natural extension of our framework is when relevant information is spread across a few input tokens rather than just one, which relates to multi-head attention. Future mathematical analyses should also consider extensions to general initialization schemes and stochastic dynamics. Our experiments (Figures 4(a), 5, and Appendix E) yield encouraging results in all these directions.
Acknowledgments
Authors thank Peter Bartlett, Linus Bleistein, Alex Damian, Spencer Frei, and Clément Mantoux for fruitful discussions and feedback. P.M. is supported by a Google PhD Fellowship.
Appendix
Organization of the Appendix.
Section A presents the main steps of Theorem 5. The intermediate results of this proof, as well as the other statements of the main text, are proven in Section B. Section C provides an expression of the risk ${\mathcal{R}}$ beyond the manifold ${\mathcal{M}}$ that extends the one provided in Lemma 6 on ${\mathcal{M}}$. Section D gives some useful technical lemmas. Experimental details and additional results are in Section E. Finally, Section F discusses additional related models.
Notation.